# Cliqueformer: Model-Based Optimization with Structured Transformers

### v1

@@ Appendix B Experimental details
We use the implementation of grudzien2024functional to generate data with latent radial-basis functions. Also, we initially wanted to use Design Bench (trabucco2022design) for experiments with practical tasks. However, at the time of this writing, the benchmark suite was suffering a data loss and was not readily available. To overcome it, we manually found the data and implemented dataset classes. TFBind-8 (trabucco2022design) could be fully downloaded since the number of possible pairs $({\mathbf{x}},{\textnormal{y}})$ is quite small. Hence, a design can be evaluated by looking up its score in the dataset. For Superconductor (hamidieh2018data), we pre-trained an XGBoost oracle on the full dataset, and trained our model and the baselines to predict the labels produced by the oracle. The proposed designs of the tested models are evaluated by calling the oracle as well. We obtained DNA Enhancers dataset from the code of uehara2024understanding, available at
@@
we additionally filter the dataset to keep only sequences featured by chromosomes from 1 to 4. We use their pre-trained oracle for generating labels and evaluation of proposed designs. Following fannjiang2020autofocused and trabucco2022design, we train our models on the portions of the datasets with values below their corresponding $80^{th}$. Upon evaluation, we obtain the ground-truth/oracle value of the proposed design y, and normalize it as
@@
and report $\bar{{\textnormal{y}}}$. ${\textnormal{y}}_{\min}$ and ${\textnormal{y}}_{\max}$ are the minimum and the maximum of the training data. This normalization scheme is different than, for example, the one in the work by trabucco2022design. We choose this scheme due to its easy interpretability—a score of $\bar{{\textnormal{y}}}>1$ implies improvement over the given dataset, which is the ultimate objective of MBO methods. However, we note that a score of less than $1$ does not imply failute of the algorithm, since we initialize our designs at a random sample from the dataset, which can be arbitrarily low-value or far from the optimum. For some functions, like in latent RBFs, the optima are very narrow spikes in a very high-dimensional space, being nearly impossible to find (see Figure 1(a)). We choose such an evaluation scheme due to its robustness that allows us to see how good ut improving any design our algorithms are overall.
@@
For baselines, we use hyper-parameters suggested by trabucco2021conservative. We decreased the hidden layer sizes (at no harm to performance) for Lat. RBF 31 and DNA Enhancers tasks where the performance was unstable with larger sizes. Also, we haven’t tuned most of the Cliqueformer’s hyper-parameters per-task. We found, however, as set of hyper-parameters that works reasonably well on all tasks.
On all tasks, we use 2 transformer blocks in both the encoder and the decoder, with transformer dimension of 64, and 2-head attention. The predictive model $f_{\theta}({\mathbf{z}})$ is a multi-layer perceptron with 2 hidden layers of dimension 256. We change it to 512 only for DNA Enhancers. The best activation function we tested was GELU (hendrycks2016gaussian), and LeakyReLU(0.3) gives similar results. We use dropout of rate 0.5 (srivastava2014dropout). In all tasks, weight of the MSE term to $\tau=$10 (recall Equation (5)). Additionally, we warm up our VIB term linearly for $1000$ steps (with maximal coefficient of 1). We train the model with AdamW (loshchilov2017fixing) with the default weight decay of Pytorch (paszke2019pytorch). We set the model learning rate to 1e-4 and the design learning rate to 3e-4 in all tasks. We train the design with AdamW with high rates of weight decay (ranging from 0.1 to 0.5).
@@
$dim({\mathbf{z}})=d_{\text{knot}}+N_{\text{clique}}\cdot(d_{\text{clique}}-d_{\text{knot}})$.

### v2

@@ Appendix D Experimental details
We use the implementation of Grudzien et al. (2024) to generate data with latent radial-basis functions. Also, we initially wanted to use Design Bench (Trabucco et al., 2022) for experiments with practical tasks. However, at the time of this writing, the benchmark suite was suffering a data loss and was not readily available. To overcome it, we manually found the data and implemented dataset classes. TFBind-8 (Trabucco et al., 2022) could be fully downloaded since the number of possible pairs $({\mathbf{x}},{\textnormal{y}})$ is quite small. Hence, a design can be evaluated by looking up its score in the dataset. For Superconductor (Hamidieh, 2018), we pre-trained an XGBoost oracle on the full dataset, and trained our model and the baselines to predict the labels produced by the oracle. The proposed designs of the tested models are evaluated by calling the oracle as well. We obtained DNA Enhancers dataset from the code of Uehara et al. (2024a), available at
@@
we additionally filter the dataset to keep only sequences featured by chromosomes from 1 to 4. We use their pre-trained oracle for generating labels and evaluation of proposed designs. Following Fannjiang & Listgarten (2020) and Trabucco et al. (2022), we train our models on the portions of the datasets with values below their corresponding $80^{th}$. Upon evaluation, we obtain the ground-truth/oracle value of the proposed design y, and normalize it as
@@
and report $\bar{{\textnormal{y}}}$. ${\textnormal{y}}_{\min}$ and ${\textnormal{y}}_{\max}$ are the minimum and the maximum of the training data. This normalization scheme is different than, for example, the one in the work by Trabucco et al. (2022). We choose this scheme due to its easy interpretability—a score of $\bar{{\textnormal{y}}}>1$ implies improvement over the given dataset, which is the ultimate objective of MBO methods. However, we note that a score of less than $1$ does not imply failute of the algorithm, since we initialize our designs at a random sample from the dataset, which can be arbitrarily low-value or far from the optimum. For some functions, like in latent RBFs, the optima are very narrow spikes in a very high-dimensional space, being nearly impossible to find (see Figure 1(a)). We choose such an evaluation scheme due to its robustness that allows us to see how good ut improving any design our algorithms are overall.
@@
For baselines, we use hyper-parameters suggested by Trabucco et al. (2021). We decreased the hidden layer sizes (at no harm to performance) for Lat. RBF 31 and DNA Enhancers tasks where the performance was unstable with larger sizes. Also, we haven’t tuned most of the Cliqueformer’s hyper-parameters per-task. We found, however, as set of hyper-parameters that works reasonably well on all tasks.
On all tasks, we use 2 transformer blocks in both the encoder and the decoder, with transformer dimension of 64, and 2-head attention. The predictive model $f_{\theta}({\mathbf{z}})$ is a multi-layer perceptron with 2 hidden layers of dimension 256. We change it to 512 only for DNA Enhancers. The best activation function we tested was GELU (Hendrycks & Gimpel, 2016), and LeakyReLU(0.3) gives similar results. We use dropout of rate 0.5 (Srivastava et al., 2014). In all tasks, weight of the MSE term to $\tau=$10 (recall Equation (3.2)). Additionally, we warm up our VIB term linearly for $1000$ steps (with maximal coefficient of 1). We train the model with AdamW (Loshchilov et al., 2017) with the default weight decay of Pytorch (Paszke et al., 2019). We set the model learning rate to 1e-4 and the design learning rate to 3e-4 in all tasks. We train the design with AdamW with high rates of weight decay (ranging from 0.1 to 0.5).
@@
$d_{z}=d_{\text{knot}}+N_{\text{clique}}\cdot(d_{\text{clique}}-d_{\text{knot}})$.